Identify the surface of the ρ = sinΦsinθ by converting them into equations in the Cartesian form. Show the complete solutions.


1
Expert's answer
2022-06-01T02:45:23-0400
"x=\rho\sin\phi\cos\theta, y=\rho\sin\phi\sin\theta, z=\rho\cos\phi"

"\rho^2=x^2+y^2+z^2"

Given

"\rho=\sin\phi\sin\theta"

Then


"\rho^2=\rho\sin\phi\sin\theta"

Substitute


"x^2+y^2+z^2=y"

"x^2+(y-\dfrac{1}{2})^2+z^2=(\dfrac{1}{2})^2"

The surface is a sphere.